OpenStudy (anonymous):

The sum of two numbers is 48. If the larger number is 6 less than twice the smaller number, find the two numbers. Write and solve a system of equations for this situation.

myininaya (myininaya):

first sentence says a number plus another number is 48. You can write that as a mathematical equation.

myininaya (myininaya):

Can you give it ago? Just translate the first sentence as a mathematical equation.

OpenStudy (anonymous):

ok so

OpenStudy (anonymous):

x+x=48

myininaya (myininaya):

we do not know the numbers are the same

myininaya (myininaya):

so x+y=48

myininaya (myininaya):

now let's assume y is the larger value

myininaya (myininaya):

second sentence says y is 6 less than twice x

myininaya (myininaya):

can you write that as a mathematical equation

OpenStudy (anonymous):

-12+x=48

myininaya (myininaya):

y is 6 less than twice x replace is with equals and n less than m means m-n and twice a number is 2 times a number

OpenStudy (anonymous):

12*4

myininaya (myininaya):

y is 6 less than twice x y = 6 less than twice x y = 6 less than 2x y=2x-6

myininaya (myininaya):

i translated each little part for you one at a time

myininaya (myininaya):

anyways first equation we had: x+y=48 second equation we had: y=2x-6

myininaya (myininaya):

insert (2x-6) for equation 1 's y

myininaya (myininaya):

replacing y with 2x-6 we have first equation now is x+2x-6=48 can you solve for x

OpenStudy (anonymous):

x=18

myininaya (myininaya):

ok can you find y

myininaya (myininaya):

recall y=2x-6

OpenStudy (anonymous):

im not sure no

myininaya (myininaya):

you remember you got x is 18?

myininaya (myininaya):

y=2(18)-6 I replaced x with 18 since x is 18

myininaya (myininaya):

2(18) means multiply 2 and 18

myininaya (myininaya):

2(18)-6 means after multiplying 2 and 18 take away 6 from that product

myininaya (myininaya):

were you able to simplify the value for y: y=2(18)-6 ?

OpenStudy (anonymous):

y=30 @myininaya

OpenStudy (anonymous):

@myininaya